OpenStudy (anonymous):

(3x) (7x^2) Please show work

OpenStudy (anonymous):

21 x^3

OpenStudy (anonymous):

just multiply

OpenStudy (anonymous):

thats all

OpenStudy (anonymous):

This is equivalent to (3x)*(7x^2) first, multiply the numbers. 3*7=21 then, multiply the unknown. x*x^2= x^3 then, put them together again. =21x^3

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you're welcome!